A 100 nC charge is placed at the origin. A -50 nC charge is placed at x= 5 cm and a 200 nC is placed at x= -10 cm.





b. What is the net electrostatic force acting on 200nC charge?


1
Expert's answer
2022-03-13T18:48:51-0400

Solution;

"F=\frac {q_1q_2}{r^2}"

"F_{200}=(\frac{100×200}{10^2}-\frac{100×50}{5^2})×8.99×10^9"

"F_{200}=0.03596N"